OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x2 + 13x + 42

OpenStudy (jhannybean):

What two numbers add to give 13, and multiply to give 42?

OpenStudy (anonymous):

is it (x+6)(x+7) ?

OpenStudy (jhannybean):

\[(x+6)(x+7) = x^2+7x+6x+42 = x^2+13x+42 ~\checkmark\]